# ChromFound: Towards A Universal Foundation Model for Single-Cell Chromatin Accessibility Data

### v2

@@ 4.2 Cell Type Annotation
Computational cell type identification is a fundamental task in single-cell omics analysis. To adapt ChromFound for this task, we introduce a feature projection head followed by MLP layers as described in Appendix F.1. Our methods are compared to three strong baseline models designed for cell type annotation of scATAC-seq, Cellcano [45], EpiAnno [6] and SANGO [86]. The variant of ChromFound trained from scratch is also included to assess the contribution of pretraining. Fig. 3 highlights that ChromFound consistently outperforms all benchmark methods, achieving the highest performance in macro F1 score and demonstrating its ability to accurately annotate diverse cell types. The confusion matrices on PBMC169K [15] for ChromFound and CellCano [45] are shown in Fig. 5. The results indicate that Cellcano struggles to correctly classify rare cell types such as natural killer cells, CD16+ monocytes, and dendritic cells, likely due to their low abundance. In contrast, ChromFound significantly improves the classification accuracy for these cell types, enabling a more precise understanding of immune cell composition and dynamics.
@@ 4.3 Cross-omics Prediction
Although recent advances have enabled the simultaneous profiling of multiple assays, most single-cell datasets remain isolated, posing significant challenges for effective multi-omics analysis. To explore the potential of ChromFound in multi-omics prediction, we utilize scATAC-seq data as the source modality and scRNA-seq data as the target modality. Implementation details are provided in Appendix A.3. For comprehensive evaluation, we benchmark on five paired scATAC-seq and scRNA-seq datasets [88, 68, 43, 4, 89] using two standard correlation metrics. Three representative baseline methods [74, 82, 72] are included with default configurations from the DANCE toolkit [16]. As shown in Table 3, ChromFound consistently outperforms all baselines across datasets in both PCC and CCC, highlighting its strong capability in cross-omics prediction. These results highlight the potential of ChromFound to facilitate biological analyses such as inferring gene-enhancer relationships and predicting transcriptional response to enhancer perturbation. The two biological applications are discussed in detail below.
@@
Enhancer elements in the human genome harbor thousands of genetic variants that regulate gene expression and contribute to the risk of common diseases. In our approach detailed in Appendix A.4, we simulate OCR knockdown and estimate the resulting changes in gene expression. The absolute change reflects the likelihood of an enhancer-gene regulatory link, while the relative change captures the direction and strength of the transcriptional response. We perform experiments on 141 K562 cells from the test split in Zhu11K [89], using the cross-omics model trained as evaluated in Table 3. Fulco4K [89] provides the ground truth of enhancer-gene links and quantitative effects through CRISPRi perturbations in K562 human erythroleukemia cells. For clarity, the evaluation focuses on the COPZ1 and HNRNPA1 genes, which are known to be involved in cancer-related processes such as stress adaptation and RNA splicing regulation [61, 7]. BABEL [74] and CMAE [82] are included as benchmark methods due to their strong performance in cross-omics prediction in Zhu11K [89]. As shown in Fig. 4, ChromFound achieves the highest AUC scores with 0.77 for COPZ1 and 0.61 for HNRNPA1, accurately identifying enhancer-gene links. Fig. 4 further shows that ChromFound better predicts transcriptional responses to enhancer perturbations. In contrast, BABEL [74] and CMAE [82] filter for the top 10,000 highly variable OCRs [16], leading to zero predictions of most enhancer perturbations. This limitation is evident in both Fig.4 and Fig.4, highlighting the advantage of ChromFound’s genome-wide modeling to support a more comprehensive understanding of regulatory mechanisms.
@@ 4.4 Ablation study
Table 4: Ablation study. Red row indicates the full model of ChromFound. The other colored rows correspond to the topics discussed in Section 4.4.
@@
As seen in rows 7-8 of Table 4, reducing the input length by half and a quarter significantly hurts performance, highlighting the importance of long-context modeling for dynamic OCR landscapes.
@@ 5 Conclusion
Appendix A Experiments Details
A.1 Cell Clustering
@@
A.2 Cell Type Annotation
@@
A.3 Cross-omics Prediction
We employ a two-layer Multilayer Perceptron (MLP) to perform the modality prediction task, with the hidden layer size set to 1024. The learning rate is configured at 1e-4, and AdamW is utilized as the optimizer. A total of 10 epochs are conducted for training. All benchmark methods are implemented with the default parameters in the DANCE package [16]. All experiments are conducted on a single machine equipped with four NVIDIA A100 GPUs. We divide the whole dataset into 80% for training, 10% for evaluation and 10% for testing.
A.4 Biological Application
@@
Appendix B Details of Datasets
@@
Appendix C Pre-processing of scATAC-seq Data
We implement a pre-process pipeline to scATAC-seq data, addressing sparsity and dimensionality to ensure high-quality cells and OCRs for model pretraining and downstream tasks. Pre-processing is performed after the train/test splits in pretraining and downstream tasks, preventing any information leakage. All datasets are saved in .h5ad format with the OCR coordinates of genome in var schema. We detail the pre-processing methods in the following.
Cell filtering: We filter cells with non-zero counts lower than 1000 or higher than 60000 for every dataset.
OCR filtering: We compute the non-zero counts of each OCR by cell types and filter OCRs with non-zero counts lower than 5% in all cell types, ensuring that long-tail cell types remain heterogeneous.
@@ Appendix C Pre-processing of scATAC-seq Data
Appendix D More related works about Mamba and hybrid architecture
Mamba-based models, such as Mamba [25] and Mamba-2 [14], replace self-attention with selective State Space Models (SSMs) [60, 27, 26], achieving linear complexity while maintaining strong performance. Recently, Mamba has been applied in biology and genomics such as RNA foundation models [18] and DNA sequence modeling [59], efficiently handling sparse and long-range dependencies in biological data.
We have noticed that some works have started to use hybrid transformer-mamba structures for modeling time series and image data. After conducting some theoretical research and analysis, we found that the existing hybrid structures are not suitable for scATAC-seq modeling. Therefore, ChromFound has been meticulously designed for scATAC-seq data, with fundamental differences from existing hybrid structures. Below, we highlight two important papers for a detailed theoretical comparison.
1. The SST [80] uses transformers and mamba to separately model short-range and long-range dependencies, then concatenates the features of short-range and long-range interactions. This resembles a parallel stacking structure. In contrast, ChromFound employs a sequential stacking approach, where data first passes through the transformer for short-range feature interactions and then moves to mamba for long-range learning.
2. MambaVision [30] integrates Vision Transformers with Mamba and is tailored for vision applications. ChromFound, however, has been designed specifically to accommodate the local strong interactions present in scATAC-seq data. We introduced a novel window partition self-attention module to better capture these interactions.
@@ Appendix D More related works about Mamba and hybrid architecture
Appendix E Discussion of comparing OCR tokenization with other peak-calling algorithms
Below, we discuss about its experimental justification and practical utility compared to approaches relying on consistent peak-calling pipelines.
1. Experimental validation and ablation study demonstrate the efficacy of genome-aware OCR tokenization: Our genome-aware tokenization framework integrates three biologically grounded components: chromosome identity, genomic coordinates, and accessibility values. By explicitly encoding positional and genomic context, ChromFound adapts to OCRs derived from heterogeneous peak-calling pipelines without requiring a unified feature space. The ablation study in 4 empirically validates this design: the inclusion of genome-aware tokenization improves clustering performance compared to a baseline lacking genomic embeddings.
2. Genome-aware tokenization reduces preprocessing complexity: A critical advantage of ChromFound lies in its ability to operate directly on precomputed count matrices (e.g., .h5ad files), eliminating dependencies on raw fragment files or reprocessing through specific peak-calling pipelines. Publicly available scATAC-seq datasets often only provide processed matrices, rendering methods reliant on unified peak sets impractical. ChromFound circumvents these limitations, ensuring immediate applicability to datasets lacking raw sequencing data. Even when fragment files are accessible, reprocessing data through unified pipelines (e.g., ENCODE) imposes prohibitive burdens. Such workflows require extensive data preprocessing, and substantial computational resources—particularly for large-scale datasets (>1 million cells). ChromFound bypasses these inefficiencies, enabling rapid integration of new data.
3. Scalability of genome-aware tokenization: The variable OCR tokenization ensures robust scalability across datasets with divergent OCR definitions. When integrating datasets with minimal OCR overlap (e.g., novel tissues or rare cell types), fixed peak sets discard biologically meaningful OCRs unique to new data. ChromFound dynamically encodes OCRs based on genomic coordinates, preserving dataset-specific regulatory elements critical for discovery-driven tasks.
Summary: ChromFound addresses two critical limitations of existing approaches: (1) dependency on raw fragment files and reprocessing workflows, and (2) inability to handle datasets with low OCR overlap. By genome-aware tokenization, our framework enables integration of heterogeneous scATAC-seq datasets without complex bioinformatic pipelines. This design aligns with the core objectives of foundation models—scalability, universality, and practical utility—as demonstrated by its state-of-the-art performance across diverse tasks.
Appendix F Analysis of Cell Type Annotation
F.1 Metric Value of Cell Type Annotation
F.2 Confusion Matrix of Cell Type Annotation
Appendix G Supplementary results of denoising batch effect
@@
We show the plots of the denoising batch effect as below. The results on To326K [68] from the Bone tissue are plotted in Fig. 6. The results on Kuppel139K [36] from the Heart tissue are plotted in Fig. 7. The results on Morabitol130K [49] from the Cortex tissue are plotted in Fig. 8. The results on PBMC169K [15] from the PBMC tissue are plotted in Fig. 9. In addition, the results of all metrics are listed in Table G.

### v3

@@ 4.2 Cell Type Annotation
Computational cell type identification is a fundamental task in single-cell omics analysis. To adapt ChromFound for this task, we introduce a feature projection head followed by MLP layers as described in Appendix B.2. Our methods are compared to three strong baseline models designed for cell type annotation of scATAC-seq, Cellcano [42], EpiAnno [7] and SANGO [82]. The variant of ChromFound trained from scratch is also included to assess the contribution of pretraining. Fig. 3 highlights that ChromFound consistently outperforms all benchmark methods, achieving the highest performance in macro F1 score and demonstrating its ability to accurately annotate diverse cell types. The confusion matrices on PBMC169K [15] for ChromFound and CellCano [42] are shown in Fig. 9. The results indicate that Cellcano struggles to correctly classify rare cell types such as natural killer cells, CD16+ monocytes, and dendritic cells, likely due to their low abundance. In contrast, ChromFound significantly improves the classification accuracy for these cell types, enabling a more precise understanding of immune cell composition and dynamics.
@@ 4.3 Cross-omics Prediction
Although recent advances have enabled the simultaneous profiling of multiple assays, most single-cell datasets remain isolated, posing significant challenges for effective multi-omics analysis. To explore the potential of ChromFound in multi-omics prediction, we utilize scATAC-seq data as the source modality and scRNA-seq data as the target modality. Implementation details are provided in Appendix B.3. For comprehensive evaluation, we benchmark on five paired scATAC-seq and scRNA-seq datasets [86, 65, 40, 5, 87] using two standard correlation metrics. Three representative baseline methods [71, 78, 69] are included with default configurations from the DANCE toolkit [16]. As shown in Table 3, ChromFound consistently outperforms all baselines across datasets in both PCC and CCC, highlighting its strong capability in cross-omics prediction. These results highlight the potential of ChromFound to facilitate biological analyses such as inferring gene-enhancer relationships and predicting transcriptional response to enhancer perturbation. The two biological applications are discussed in detail below.
@@
Enhancer elements in the human genome harbor thousands of genetic variants that regulate gene expression and contribute to the risk of common diseases [85]. In our approach detailed in Appendix B.4, we simulate OCR knockdown and estimate the resulting changes in gene expression. The absolute change reflects the likelihood of an enhancer-gene regulatory link, while the relative change captures the direction and strength of the transcriptional response. We perform experiments on 141 K562 cells from the test split in Zhu11K [87], using the cross-omics model trained as evaluated in Table 3. Fulco4K [87] provides the ground truth of enhancer-gene links and quantitative effects through CRISPRi perturbations in K562 human erythroleukemia cells. For clarity, the evaluation focuses on the COPZ1 and HNRNPA1 genes, which are involved in cancer-related processes such as stress adaptation and RNA splicing regulation [58, 8]. BABEL [71] and CMAE [78] are included as benchmark methods due to their strong performance in cross-omics prediction in Zhu11K [87]. As shown in Fig. 4, ChromFound achieves the highest AUC scores with 0.77 for COPZ1 and 0.61 for HNRNPA1, accurately identifying enhancer-gene links. Fig. 4 further shows that ChromFound better predicts transcriptional responses to enhancer perturbations. In contrast, BABEL [71] and CMAE [78] filter for the top 10,000 highly variable OCRs [16], leading to zero predictions of most enhancer perturbations. This limitation is evident in both Fig.4 and Fig.4, highlighting the advantage of ChromFound’s genome-wide modeling to support a more comprehensive understanding of regulatory mechanisms.
@@ 4.4 Ablation study
As seen in rows 7-8 of Table 4, reducing the input length by half and by a quarter significantly hurts performance, highlighting the importance of long-context modeling for dynamic OCR landscapes.
Table 4: Ablation study. Red row indicates the full model of ChromFound. The other colored rows correspond to the topics discussed in Section $4.$
@@ 5 Conclusion
6 Acknowledgments
The authors gratefully acknowledge the following funding sources: AI for Science Program, Shanghai Municipal Commission of Economy and Information, National Natural Science Foundation of China (Grant No. 82394432, 92249302), Shanghai Municipal Science and Technology Major Project (2023SHZDZX02), the King Abdullah University of Science and Technology (KAUST) Office of Research Administration (ORA) under Award No REI/1/5234-01-01, REI/1/5414-01-01, REI/1/5289-01-01, REI/1/5404-01-01, REI/1/5992-01-01, URF/1/4663-01-01, Center of Excellence for Smart Health (KCSH) under award number 5932, and Center of Excellence on Generative AI under award number 5940. The computations in this research were performed using the Computing for the Future at Fudan (CFFF) platform of Fudan University.
Appendix A Discussion of comparing OCR tokenization with other peak-calling algorithms
Traditional scATAC-seq pipelines often rely on a fixed OCR vocabulary (e.g., from a reference atlas) to harmonize datasets, which ensures alignment consistency but introduces two major limitations: (1) a loss of OCR length information, which encodes accessibility strength and resolution, and (2) reduced generalizability, as tissue-specific or novel OCRs are often excluded. ChromFound addresses these issues through a genomic-coordinate–based tokenization strategy that encodes each OCR by its actual genomic start and end positions using sinusoidal positional embeddings (Section 3.1). This position-aware representation naturally accommodates variable-length OCRs and preserves their spatial relationships, enabling the model to capture regulatory heterogeneity across tissues.
We perform an experiment to directly compare ChromFound with VAE-based models trained on reference-peak-aligned data. Specifically, we adopt the cPeaks reference set [45], which defines 1,657,194 peaks across the human genome. We use a 20,000-cell PBMC subset (same as in Table 4, Row 6) as the training data. After mapping peaks to the cPeaks reference and filtering out peaks and cells with all-zero values, we obtain a training set of 306,784 peaks and approximately 18,000 cells.
We train four VAE-based baselines [73, 72, 1, 44] on the aligned data and compare their performance on zero-shot cell clustering (Table 1) against ChromFound trained on the same aligned data. Evaluation metrics and datasets follow Table1, all aligned to 306,784 peaks. Results are summarized in Table 5.
This comparison yields two key observations. First, reference-based alignment substantially limits cross-tissue generalization: only 30% of peaks in retina, 35% in cortex, and 70% in heart overlap with the cPeaks-aligned PBMC training peaks, leading to markedly lower performance of the baseline VAEs on unseen tissues. Second, ChromFound, though robust even when trained on the reference-aligned data, performs best when using its native OCRs, suggesting that fixed-peak harmonization fails to capture novel or shifted OCRs arising from tissue heterogeneity, disease-specific variation, or batch effects.
Table 5: Results of VAE-based models with cPeaks-aligned training against ChromFound. Result style: best, second best
Dataset Model VAE cPeaks ARI($\uparrow$) FMI($\uparrow$) NMI($\uparrow$) AMI($\uparrow$)
SCALE [73] 0.2932 0.5050 0.4127 0.4081
SCALEX [72] 0.4230 0.6078 0.5543 0.5513
peakVI [1] 0.3690 0.5643 0.5023 0.4983
poissonVI [44] 0.3364 0.5408 0.4636 0.4594
ChromFound 0.5858 0.7244 0.7519 0.7499
Cortex(batch 1) ChromFound 0.6475 0.7503 0.7701 0.7689
SCALE [73] 0.2484 0.4052 0.3657 0.3609
SCALEX [72] 0.4481 0.5679 0.5742 0.5710
peakVI [1] 0.3610 0.5035 0.4876 0.4835
poissonVI [44] 0.2820 0.4374 0.4104 0.4058
ChromFound 0.5751 0.6720 0.7145 0.7123
Cortex(batch 2) ChromFound 0.6092 0.7015 0.7276 0.7259
SCALE [73] 0.3543 0.4817 0.5228 0.5195
SCALEX [72] 0.2795 0.4150 0.4787 0.4752
peakVI [1] 0.3023 0.4437 0.4816 0.4779
poissonVI [44] 0.3778 0.5025 0.5446 0.5414
ChromFound 0.5400 0.6347 0.6588 0.6567
Heart(av 3) ChromFound 0.5642 0.6601 0.6981 0.6953
SCALE [73] 0.4051 0.6506 0.4461 0.4453
SCALEX [72] 0.3013 0.5685 0.4199 0.4191
peakVI [1] 0.3966 0.6080 0.5018 0.5011
poissonVI [44] 0.3939 0.5997 0.5010 0.5003
ChromFound 0.6150 0.7695 0.6968 0.6963
Heart(av 10) ChromFound 0.6432 0.7869 0.7220 0.7223
SCALE [73] 0.3065 0.5771 0.4447 0.4440
SCALEX [72] 0.2430 0.5149 0.4127 0.4117
peakVI [1] 0.2514 0.5260 0.4651 0.4644
poissonVI [44] 0.2347 0.5122 0.4063 0.4055
ChromFound 0.4888 0.7081 0.7241 0.7238
Retina(D026_13) ChromFound 0.5802 0.7549 0.7410 0.7403
SCALE [73] 0.3143 0.5092 0.4220 0.4210
SCALEX [72] 0.2490 0.4542 0.3920 0.3909
peakVI [1] 0.2664 0.4693 0.4745 0.4736
poissonVI [44] 0.2396 0.4473 0.3937 0.3926
ChromFound 0.5368 0.6839 0.7664 0.7660
Retina(D19D008) ChromFound 0.5956 0.7202 0.7832 0.7868
SCALE [73] 0.5769 0.6756 0.7066 0.7057
SCALEX [72] 0.5718 0.6514 0.7205 0.7196
peakVI [1] 0.5306 0.6249 0.7057 0.7046
poissonVI [44] 0.5593 0.6479 0.6742 0.6732
ChromFound 0.5999 0.6827 0.7434 0.7424
PBMC(VIB_10xv1_1) ChromFound 0.6142 0.6995 0.7354 0.7345
SCALE [73] 0.4031 0.5797 0.5380 0.5374
SCALEX [72] 0.3326 0.5430 0.5465 0.5458
peakVI [1] 0.3830 0.5804 0.5737 0.5731
poissonVI [44] 0.3959 0.5917 0.5509 0.5503
ChromFound 0.4347 0.6237 0.5867 0.5861
PBMC(BIO_ddseq_1) ChromFound 0.4579 0.6421 0.5902 0.5910
Appendix B Experiments details
B.1 Cell clustering
@@
B.2 Cell type annotation
@@
B.3 Cross-omics prediction
We employ a two-layer Multilayer Perceptron (MLP) to perform the modality prediction task, with the hidden layer size set to 1024. The learning rate is configured at 1e-4, and AdamW is utilized as the optimizer. A total of 10 epochs are conducted for training. All benchmark methods are implemented with the default parameters in the DANCE package [16]. All experiments are conducted on a single machine equipped with four NVIDIA A100 GPUs. We divide the whole dataset into 80% for training, 10% for evaluation, and 10% for testing.
B.4 Biological application
@@
Appendix C Details of datasets
@@
Appendix D Pre-processing of scATAC-seq data
We develop a preprocessing pipeline for scATAC-seq data to address sparsity and high dimensionality, ensuring high-quality cells and OCRs for model pretraining and downstream analyses. Preprocessing is applied after train/test splitting to avoid information leakage, and all datasets are stored in .h5ad format with genomic OCR coordinates in the var schema.
Quality Control: We filter out cells with non-zero counts below 1,000 or above 60,000 and remove OCRs with non-zero counts present in less than 5% of all cell types to preserve heterogeneity among rare populations. For unlabeled datasets, pseudo labels are inferred using TF-IDF followed by LSI for OCR filtering.
Cell aggregation: To mitigate sparsity, we aggregate neighboring cells identified through Latent Semantic Indexing (LSI) and cosine similarity. LSI extracts $n$ components after removing technical noise (the first component), and cosine similarity determines the 10 nearest neighbors for aggregation.
@@ Appendix D Pre-processing of scATAC-seq data
Appendix E Supplementary results of ablation study
E.1 Hyperparameter sensitivity
We evaluate the hyperparameter sensitivity on the cell clustering task using PBMC169K (batch VIB_10xv1_1) dataset.
Positional embedding temperature: This parameter controls the frequency of sinusoidal position encodings and can be viewed as a proxy for "genomic resolution". As shown in Table 7, performance slightly drops when temp is within 10e3 to 10e5, and degrades more substantially when temp is too large, likely due to loss of relative position sensitivity.
Table 7: Results of hyperparameter temptemp sensitivity experiment.
temptemp ARI($\uparrow$) NMI($\uparrow$) AMI($\uparrow$) FMI($\uparrow$)
1000 0.6535 0.7764 0.7755 0.7259
10000 0.6512 0.7709 0.7701 0.7243
100000 (Ours) 0.6953 0.7860 0.7852 0.7601
1000000 0.6036 0.7344 0.7334 0.6860
10000000 0.5852 0.7269 0.7258 0.6714
Mamba projection dimension: This parameter controls the compression within the Mamba block. As shown in Table 8, $D_{low}$ = 32 achieves a strong balance between performance and efficiency. Our choice of 32 is primarily motivated by the trade-off between performance and computational efficiency. Larger values lead to marginal gains but significantly increase FLOPs and memory, with $D_{low}$ = 128 exceeding the memory limits on A100 80G GPUs.
Table 8: Results of Mamba projection dimension $D_{low}$ sensitivity experiment.
$D_{low}$ #Parameter FLOPs (10e11) Inference Speed (s) VRAM (GB) ARI($\uparrow$)
16 422,593 7.41 3.4027 48.8 0.6158
32 (Ours) 450,305 7.89 3.4624 60.7 0.6953
64 518,785 9.09 4.0053 72.3 0.6927
128 707,969 12.4 OOM OOM OOM
E.2 Comparison on Transformer-only architectures
Training a vanilla Transformer on scATAC-seq inputs containing nearly one million peaks is practically infeasible due to its quadratic scaling in memory and computation. To approximate a pure Transformer baseline under practical settings, we train two representative single-cell foundation models, Geneformer and scGPT, on the same pretraining corpus with increasing peak lengths (4k–32k) and the same hyperparameters as WPSA. The results of comparison on the cell clustering task using the PBMC169K (batch VIB_10xv1_1) dataset are detailed in Table 9.
Table 9: Results of comparison on Transformer-only architectures.
Method OCRs Length ARI($\uparrow$)
Geneformer 4096 0.0451
16384 0.0457
32768 0.0460
scGPT 4096 0.3075
16384 0.3774
32768 0.3868
ChromFound 440,000 0.6953
Appendix F Supplementary results of denoising batch effect
@@
We show the plots of the denoising batch effect as below. The results on To326K [65] from the Bone tissue are plotted in Fig. 5. The results on Kuppel139K [33] from the Heart tissue are plotted in Fig. 6. The results on Morabitol130K [48] from the Cortex tissue are plotted in Fig. 7. The results on PBMC169K [15] from the PBMC tissue are plotted in Fig. 8. In addition, the results of all metrics are listed in Table 10.
@@
Appendix G Analysis of Cell Type Annotation
G.1 Metric Value of Cell Type Annotation
G.2 Confusion Matrix of Cell Type Annotation
@@ H.1.3 CASTLE
H.1.5 PeakVI
PeakVI [1] is a variational autoencoder (VAE)–based probabilistic model for single-cell chromatin accessibility data. It captures both biological and technical variation through a hierarchical generative process, enabling effective batch correction and denoising. The learned latent representation supports diverse downstream tasks such as clustering and differential accessibility analysis, providing a strong probabilistic baseline for scATAC-seq modeling.
H.1.6 PoissonVI
PoissonVI [44] extends the VAE framework by modeling raw accessibility counts with a Poisson likelihood instead of binarized inputs. This design preserves quantitative accessibility information and improves sensitivity to subtle regulatory differences. PoissonVI achieves robust performance across clustering and integration tasks while maintaining a clear probabilistic interpretation.
H.1.7 SnapATAC2
SnapATAC2 [84] is a scalable framework for large-scale scATAC-seq analysis. It represents each cell by a high-dimensional accessibility profile and applies diffusion- and spectral-based dimensionality reduction to reveal chromatin structure. With optimized memory management, parallel I/O, and GPU acceleration, SnapATAC2 efficiently handles datasets with millions of cells and serves as a strong non–deep learning baseline.
H.1.8 Signac
Signac [59] is an R package built on the Seurat ecosystem for single-cell chromatin accessibility analysis. It integrates preprocessing, feature selection, dimensionality reduction, and visualization within a unified workflow, and supports multimodal integration with scRNA-seq data. Owing to its usability and reproducibility, Signac remains a widely adopted toolkit for exploratory and integrative single-cell epigenomic analysis.